if x=9 and y=6 evaluate the expression
8-y
x-5

Answer by pmesler(52)

All you need to do for this problem is plug in the numbers that you know into the given equations.
If x = 9, then x-5 = 4 because 9-5 = 4.
If y = 6, then 8-y = 2 because 8-6 = 2.